# Optimal Piecewise-based Mechanism for Collecting Bounded Numerical Data under Local Differential Privacy

### v1

@@ 8. Related Work
For bounded data domains, the basic idea is to sample the output from a carefully designed distribution on the bounded domain. Duchi et al. randomize any data in $[-1,1)$ to two discrete values $y\in\{-C_{\varepsilon},C_{\varepsilon}\}$ (Duchi et al., 2016), where $C_{\varepsilon}$ is a constant depending on the privacy level $\varepsilon$. This binary-output mechanism exhibits a large randomization error as the output space is too coarse.  PM (Wang et al., 2019) extends the binary output to a continuous output in $[-C_{\varepsilon},C_{\varepsilon})$. It designs a piecewise-based mechanism to sample the output, which uses different sampling intervals for different $x$, achieving a lower randomization error. Both PM and later SW (Li et al., 2020) have shown the data utility advantage of TPM in numerical data collection with bounded domains under LDP. PTT (Ma et al., 2024) discusses the optimality of TPM. It shows that there exist TPM instantiations that yield optimal data utility, but it does not provide the closed-form mechanism.  TPM is a special case of GPM using $3$-piece distributions and a specific form for each piece. Meanwhile, existing TPM instantiations are designed for specific error metrics and only classical domains.
Applications of TPM. A natural application of TPM is in high-dimensional numerical data. This includes scenarios where the sensitive data may be a high-dimensional vector (Wang et al., 2019; Zhang et al., 2023), an infinite data stream (Ren et al., 2022), or matrixes (Wang et al., 2023).  Another application area is federated learning. TPM ensures LDP in the $[0,1)$ domain, which is commonly used as the normalized domain in model training. TPM avoids the data clipping required by the noise-adding mechanisms (Abadi et al., 2016; Liu et al., 2023; Fu et al., 2024).  Another typical application is in sensors, where the data is bounded by the sensor’s physical nature. This paper can replace the existing TPM to achieve better utility.
@@ Appendix A Proofs
Denote $f_{i}(x)\coloneq\int_{l_{i}}^{r_{i}}\mathcal{L}(y,x)\mathrm{d}y$, where $\mathcal{L}(y,x)=|y-x|^{p}$. First, we prove that $f_{i}(x)$ is a convex function w.r.t. $x$.  Based on the relationship between $x$ and $[l_{i},r_{i})$, the value of $x$ is split into three cases: (i) $x\in[a,l_{i})$, (ii) $x\in[l_{i},r_{i})$, and (iii) $x\in[r_{i},b)$. We prove the second derivative of $f_{i}(x)$ w.r.t. $x$ is non-negative in each case, thus $f_{i}(x)$ is convex.
@@
This lemma means that we can determine the optimal $m$-piecewise mechanism for $m=1,2,\ldots$, until the optimal $(m+1)$-piecewise mechanism is identical to the optimal $m$-piecewise mechanism for all $x$ and $\varepsilon$. This statement can be empirically validated by attempting to find counterexamples using larger $m$ than the optimal $m$. The source code of our framework provides this validation.
@@
Remark: This proof is the same as finding the optimal $3$-piecewise distribution in domain $[0,1)$. The source code of our framework provides this validation.
@@ Appendix B Complementary Materials
In this paper, the optimality of GPM is defined with respect to: (i) the worst-case $L_{p}$-similar error metric, (ii) bounded numerical domains $\mathcal{D}\to\tilde{\mathcal{D}}$ and mechanisms based on piecewise distributions, (iii) minimization of error value (not asymptotic or order-of-magnitude optimality), and (iv) without post-processing. $L_{p}$-similar error metrics are natural choices for evaluating data utility (Hardt and Talwar, 2010; Geng and Viswanath, 2014; Wang et al., 2019). Bounded numerical domains are common in real-world applications. Focusing on error values allows for more precise comparisons between different mechanisms. By excluding post-processing, we can analyze the optimality of the mechanism itself, providing a more fundamental understanding than considering the mechanism combined with specific post-processing.
Other types of optimality have been explored in the literature, particularly for variants of Laplace mechanisms. The staircase mechanism (Geng and Viswanath, 2014) adopts the same utility model without prior knowledge or post-processing as this paper. It claims optimality under specific assumptions, one of which is that a staircase (piecewise) distribution can achieve the optimal error. The mechanism demonstrates better $L_{1}$-error performance than the Laplace mechanism on $\tilde{\mathcal{D}}=(-\infty,\infty)$, and its asymptotic optimality has been formally proven.  Universal optimality is another type of optimality, defined from the perspective of a user’s prior knowledge and post-processing ability (Ghosh et al., 2009). In this utility model, the user observes the output of the mechanism and selects another value based on the output and their prior knowledge, i.e. under a Bayesian utility framework. Formally, if the user’s prior is denoted as $p_{i}$ on the data domain $i\in N$ (i.e. a discrete domain) and the user’s post-processing is represented as a remap $z_{i,j}$ that reinterprets the output of the mechanism (on the sensitive value $i$) to $j$, then the utility model is
@@
This utility model incorporates the user’s prior knowledge and post-processing ability. A mechanism is called universally optimal if, for any prior $p_{i}$, there exists an optimal remap $z_{i,j}$. Under this utility model, it was proven that the truncated geometric mechanism (a discretized version of the Laplace mechanism) can achieve universal optimality for count queries111111 This is in the centralized DP setting, where the data curator holds the dataset and uses one mechanism. and a legal error metric $\mathcal{L}(i,j)$. Such universal optimality was shown to be unachievable for more complex queries (Brenner and Nissim, 2014). Under the same utility model, the universal optimality was extended to the truncated Laplace mechanism for a bounded numerical domain $\mathcal{D}=[0,1]$ by approximating the geometric mechanism with the Laplace mechanism and post-processing (Fernandes et al., 2021).
These results do not hold in our utility model, i.e. utility model without prior and post-processing. Figure 13 has shown that OGPM generally has a smaller error than the truncated Laplace mechanism, indicating the sub-optimality of the truncated Laplace mechanism in the absence of prior and post-processing.
@@
Direction 1: If we can further move the green piece in Figure 19(b) “into” the red central piece while keeping the probabilities of the red and blue pieces unchanged, i.e. transform it into a $3$-staircase distribution while ensuring a decrease in the error, then we can prove that the optimal $m$ is $3$. However, this is challenging, as it breaks the probability normalization constraint, requiring adjustments to the probabilities of each piece to satisfy the $\varepsilon$-LDP constraint. The difficulty lies in ensuring that these adjustments will indeed decrease the error.
Direction 2: Another approach is to formulate the problems for $3$-staircase and $4$-staircase distributions as two constrained optimization problems. The goal would be to prove that the optimal error of the $3$-staircase distribution is equivalent to that of the $4$-staircase distribution. Ideally, these two multi-variable optimization problems could be solved analytically, resulting in two closed-form error expressions w.r.t. $x$ and $\varepsilon$, thereby completing the proof for any $x$ and $\varepsilon$. This direction aligns with our framework. However, the challenge lies in the complexity of solving such multi-variable optimization problems analytically. This is why we rely on an off-the-shelf optimization solver, which, while effective, only provides numerical solutions for specific $x$ and $\varepsilon$ values.
@@
$MSE(\mathcal{M}(0),0)=\frac{p_{\varepsilon}}{3}\left(8C^{3}\right)+\frac{p_{\varepsilon}}{3\exp(\varepsilon)}\left(1-8C^{3}\right).$
Setting $\varepsilon=1$ results in $MSE(\mathcal{M}(0),0)=0.22$. As a comparison, SW (Li et al., 2020), which also designed for $\mathcal{D}=[0,1)$, has an MSE of $0.29$ at $x=0$.
@@ B.6. Optimal Assignment of Privacy Parameter and an Example (Section 6)
The objective function to minimize the total error in 2D polar coordinates is
@@
the classical domain $[0,1)$ is achieved at $x=0$. Therefore, $Err_{\text{1,wor}}$ equals to $MSE(\mathcal{M}(0),0)$ calculated before. For the circular domain $[0,2\pi)$, the worst-case error $Err_{\text{2,wor}}$ is stated in Theorem 4.4.  If $\varepsilon=1+2\pi$ and we assign $\varepsilon_{1}$ to the classical domain and $\varepsilon_{2}=\varepsilon-\varepsilon_{1}$ to the circular domain, then the total error is plotted in the right figure. The optimal assignment is $\varepsilon_{1}=1.32$ and $\varepsilon_{2}=5.69$.
@@
The B-Laplace mechanism redefines a Laplace-shaped distribution on a bounded domain as the perturbation mechanism. For the domain $\mathcal{D}\to\tilde{\mathcal{D}}=[0,1)\to[0,1)$, the B-Laplace mechanism is defined as follows:

### v2

@@ 8. Related Work
For bounded data domains, the basic idea is to sample the output from a carefully designed distribution on the bounded domain. Duchi et al. randomize any data in $[-1,1]$ to two discrete values $y\in\{-C_{\varepsilon},C_{\varepsilon}\}$ (Duchi et al., 2016), where $C_{\varepsilon}$ is a constant depending on the privacy level $\varepsilon$. This binary-output mechanism exhibits a large randomization error as the output space is too coarse.  PM (Wang et al., 2019) extends the binary output to a continuous output in $[-C_{\varepsilon},C_{\varepsilon}]$. It designs a piecewise-based mechanism to sample the output, which uses different sampling intervals for different $x$, achieving a lower randomization error. Both PM and later SW (Li et al., 2020) have shown the data utility advantage of TPM in numerical data collection with bounded domains under LDP. PTT (Ma et al., 2024) discusses the optimality of TPM. It shows that there exist TPM instantiations that yield optimal data utility, but it does not provide the closed-form mechanism.  TPM is a special case of GPM using $3$-piece distributions and a specific form for each piece. Meanwhile, existing TPM instantiations are designed for specific error metrics and only classical domains.
Applications of TPM. A natural application of TPM is in high-dimensional numerical data. This includes scenarios where the sensitive data may be a high-dimensional vector (Wang et al., 2019; Zhang et al., 2023), an infinite data stream (Ren et al., 2022), or matrixes (Wang et al., 2023).  Another application area is federated learning. TPM ensures LDP in the $[0,1]$ domain, which is commonly used as the normalized domain in model training. TPM avoids the data clipping required by noise-adding mechanisms (Abadi et al., 2016; Liu et al., 2023; Fu et al., 2024).  Another typical application is in sensors, where the data is bounded by the sensor’s physical nature. This paper can replace the existing TPM to achieve better utility.
@@ 9. Conclusions
Acknowledgements.
We thank the anonymous reviewers and the revision editor for their valuable feedback and guidance, which significantly improved this paper. We also acknowledge the use of GPT-4 for language refinement in this paper.  This research is supported in part by the U.S. National Science Foundation under grants CNS-2245689 (CRII) and DGE-2336252, as well as by the 2022 Meta Research Award for Privacy-Enhancing Technologies.
@@ Appendix A Proofs
Denote $f_{i}(x)\coloneq\int_{l_{i}}^{r_{i}}\mathcal{L}(y,x)\mathrm{d}y$, where $\mathcal{L}(y,x)=|y-x|^{p}$. First, we prove that $f_{i}(x)$ is a convex function w.r.t. $x$.  Specifically, based on the relationship between $x$ and $[l_{i},r_{i})$, the value of $x$ is split into three cases: (i) $x\in[a,l_{i})$, (ii) $x\in[l_{i},r_{i})$, and (iii) $x\in[r_{i},b)$. We prove the second derivative of $f_{i}(x)$ w.r.t. $x$ is non-negative in each case, thus $f_{i}(x)$ is convex.
@@
This lemma means that we can determine the optimal $m$-piecewise mechanism for $m=1,2,\ldots$, until the optimal $(m+1)$-piecewise mechanism is identical to the optimal $m$-piecewise mechanism for all $x$ and $\varepsilon$. This statement can be empirically validated by attempting to find counterexamples using larger $m$ than the optimal $m$. The source code of our framework provides scripts and results to empirically validate this lemma.
@@
Remark: This proof is the same as finding the optimal $3$-piecewise distribution in domain $[0,1)$. The source code of our framework provides the validation.
@@ Appendix B Complementary Materials
In this paper, the optimality of GPM is defined with respect to: (i) the worst-case $L_{p}$-similar error metric, (ii) bounded numerical domains $\mathcal{D}\to\tilde{\mathcal{D}}$ and mechanisms based on piecewise distributions, (iii) minimization of error value (not asymptotic or order-of-magnitude optimality), and (iv) without post-processing. $L_{p}$-similar error metrics are natural choices for evaluating data utility (Hardt and Talwar, 2010; Geng and Viswanath, 2014; Wang et al., 2019). Bounded numerical domains are common in real-world applications. Focusing on error values allows for more precise comparisons between different mechanisms. By excluding post-processing, we can analyze the optimality of the mechanism itself, which provides a more fundamental understanding than considering the mechanism combined with a specific post-processing.
Other types of optimality have been explored in the literature, particularly for variants of Laplace mechanisms. The staircase mechanism (Geng and Viswanath, 2014) adopts the same utility model without prior knowledge or post-processing as this paper. It claims optimality under specific assumptions, one of which is that a staircase (piecewise) distribution can achieve the optimal error. The mechanism demonstrates better $L_{1}$-error performance than the Laplace mechanism on $\tilde{\mathcal{D}}=(-\infty,\infty)$, and its asymptotic optimality has been formally proven.  Universal optimality is another type of optimality, defined from the perspective of a user’s prior knowledge and post-processing ability (Ghosh et al., 2009). In this utility model, the user observes the output of the mechanism and selects another value based on the output and their prior knowledge, i.e. under a Bayesian utility framework. Formally, if the user’s prior is denoted as $p_{i}$ on the data domain $i\in N$ (i.e. a discrete domain) and the user’s post-processing is represented as a remap $z_{i,j}$ that reinterprets the output of the mechanism (on the sensitive value $i$) to $j$, then the utility model is defined as
@@
This utility model incorporates the user’s prior knowledge and post-processing ability. A mechanism is called universally optimal if, for any prior $p_{i}$, there exists an optimal remap $z_{i,j}$. Under this utility model, it was proven that the truncated geometric mechanism (a discretized version of the Laplace mechanism) can achieve universal optimality for count queries101010 This is in the centralized DP setting, where the data curator holds the dataset and uses one mechanism. and a legal error metric $\mathcal{L}(i,j)$. Such universal optimality was shown to be unachievable for more complex queries (Brenner and Nissim, 2014). Under the same utility model, the universal optimality was extended to the truncated Laplace mechanism for a bounded numerical domain $\mathcal{D}=[0,1]$ by approximating the geometric mechanism with the Laplace mechanism and post-processing (Fernandes et al., 2021).
These results do not hold in our utility model, i.e. utility model without prior and post-processing. Figure 13 has shown that OGPM generally has a smaller error than the truncated Laplace mechanism, especially when the privacy parameter $\varepsilon$ is not small, indicating the sub-optimality of the truncated Laplace mechanism in the absence of using prior and post-processing.
@@
Direction 1: If we can further move the green piece in Figure 19(b) “into” the red central piece while keeping the probabilities of the red and blue pieces unchanged, i.e. transform it into a $3$-staircase distribution while ensuring a decrease in the error, then we can prove that the optimal $m$ is $3$. However, this is challenging, as it breaks the probability normalization constraint (i.e. the sum of probabilities is no longer $1$), requiring adjustments to the probabilities of each piece to satisfy the $\varepsilon$-LDP constraint. The difficulty lies in ensuring that these adjustments will indeed decrease the error.
Direction 2: Another approach is to formulate the problems for $3$-staircase and $4$-staircase distributions as two constrained optimization problems. The goal would be to prove that the optimal error of the $3$-staircase distribution is equivalent to that of the $4$-staircase distribution. Ideally, these two multi-variable optimization problems could be solved analytically, resulting in two closed-form error expressions w.r.t. $x$ and $\varepsilon$, thereby completing the proof for any $x$ and $\varepsilon$ by showing that the two expressions are equal. This direction aligns with our framework. However, the challenge lies in the complexity of solving such multi-variable optimization problems analytically. This is why we rely on an off-the-shelf optimization solver, which, while effective, only provides numerical solutions for specific $x$ and $\varepsilon$ values.
@@ B.5. MSE of the Optimal GPM (Section 4)
$\mathrm{MSE}[\mathcal{M}(0)]=\frac{p_{\varepsilon}}{3}\left(8C^{3}\right)+\frac{p_{\varepsilon}}{3\exp(\varepsilon)}\left(1-8C^{3}\right).$
Setting $\varepsilon=1$ results in $\mathrm{MSE}[\mathcal{M}(0)]=0.22$ of OGPM. As a comparison, SW (Li et al., 2020), which also designed for $\mathcal{D}=[0,1)$, has an MSE of $0.29$ at $x=0$.
@@ B.6. Optimal Assignment of Privacy Parameter and an Example (Section 6)
The objective function to minimize the given total error in 2D polar coordinates is
@@
the classical domain $[0,1)$ is achieved at $x=0$. Therefore, $Err_{\text{1,wor}}$ equals to $\mathrm{MSE}[\mathcal{M}(0)]$ calculated before. For the circular domain $[0,2\pi)$, the worst-case error $Err_{\text{2,wor}}$ is stated in Theorem 4.4.  If $\varepsilon=1+2\pi$ and we assign $\varepsilon_{1}$ to the classical domain and $\varepsilon_{2}=\varepsilon-\varepsilon_{1}$ to the circular domain, then the total error is plotted in the right figure. In this figure, the optimal assignment is $\varepsilon_{1}=1.32$ and $\varepsilon_{2}=5.69$.
@@
The B-Laplace mechanism redefines a Laplace-shaped distribution on a bounded domain as the perturbation mechanism. For the data domain $\mathcal{D}\to\tilde{\mathcal{D}}=[0,1)\to[0,1)$, the B-Laplace mechanism is defined as follows: